A Concentration cell was constructed by immersing two silver electrodes in 0.05M and 1M AgNO3 solution. Write the cell representation , cell reactions and calculate the emf of the cell.
1
Expert's answer
2019-09-11T05:23:01-0400

Solution.

The cell representation:

"Ag^{1+}|Ag^0||Ag^{1+}|Ag^0"

The cell reactions:

But first, you need to determine the cathode and anode. In the concentration element, the cathode is the electrode with the solution, the concentration of the solute of which is greater. And the electrode with the solution, the concentration of the dissolved substance of which is less, is the anode.

"A(-): Ag^0 = Ag^{1+} + 1e"

"C(+): Ag^{1+} + 1e = Ag^0"

The emf of the cell:

To find the EMF of a concentration element, you must use the following formula:

"E = \frac{RT}{zF} \times ln(\frac{c1}{c2})"

c1>c2

E = 0.08 V

Answer:

The cell representation:

"Ag^{1+}|Ag^0||Ag^{1+}|Ag^0"

The cell reactions:

"A(-): Ag^0 = Ag^{1+} + 1e"

"C(+): Ag^{1+} + 1e = Ag^0"

The emf of the cell:

E = 0.08 V